OpenStudy (anonymous):

Tell whether the following system has one solution, infinitely many solutions, or no solution. x/4 + y/9 = 27 3x/2 + 2y/3 = 19

OpenStudy (anonymous):

@primeralph @nonamenever @skullpatrol @chetan552 @MrNebula @amistre64

OpenStudy (mrnebula):

u like my help?

OpenStudy (anonymous):

Yes, please.

OpenStudy (mrnebula):

look like a code for a safe

OpenStudy (anonymous):

It says it has no solution but I need an explanation..?

OpenStudy (anonymous):

\[27 \neq 19\] Thus it is no solution.

OpenStudy (amistre64):

multiply the top and bottom by a common factor to get rid of the fractions

OpenStudy (amistre64):

9,4,2,3 all seem to have 36 in common 36[x/4 + y/9 = 27] 36[3x/2 + 2y/3 = 19] 9x + 4y = 27(36) 3(18)x + 2(12)y = 19(36) 9x + 4y = 27(36) 3(2*9)x + 2(3*4)y = 19(36) looks like we can factor out a 6 to compare with 9x + 4y = 27(36) 9x + 4y = 19(6)

OpenStudy (anonymous):

For more detail explanation follow this link... http://www.algebra.com/algebra/homework/coordinate/Linear-systems.faq.question.320609.html

OpenStudy (amistre64):

since both left had sides are exactly the same, the right hand sides would have to be identical as well does 27(36) = 19(6) ??

OpenStudy (anonymous):

No it doesn't, I see what you mean.

OpenStudy (amistre64):

this of course is not the only way toapproach this, it was just what was on the top of my head :)

OpenStudy (anonymous):

Well you explained it very well, thank you so much c:

OpenStudy (amistre64):

youre welcome, and good luck :)

OpenStudy (anonymous):

Thanks c:

OpenStudy (anonymous):

@LanguageEnthusiast @chetan552 Thank you as well c: